A woman of weight 300N wants ro carry a load weight at 150N to the top of hill 30m high. to reach the hill top, she walk along a path of 450m. How much work has she done to carry herself and the load to the hill.
1
Expert's answer
2019-08-01T09:55:30-0400

The work done by the woman to carry herself to the hill is 300N*30m = 9000N*m = 9000J, the work done to carry the load is 150N*30m = 4500J. The total work is 9000J + 4500j = 13500J.

Answer: 13500J.